OpenStudy (anonymous):

Given the functions f(x) = 4x2 − 1, g(x) = x2 − 8x + 5, and h(x) = –3x2 − 12x + 1, rank them from least to greatest based on their axis of symmetry.

OpenStudy (anonymous):

@robotoby

OpenStudy (anonymous):

There are two different formulas that you can use to find the axis of symmetry. One formula works when the parabola's equation is in vertex form and the other works when the parabola's equation is in standard form. If your equation is in vertex form, then the axis of is: x= h in the general vertex form equation y = (x-h)2 + k If your equation is in standard form, then the formula for the axis of symmetry is: x = -b/2a from the general standard form equation y = ax2+bx + c